OpenStudy (anonymous):

A spike of a wheel reaches from the center of the wheel to its rim. If the circumference of the rim of the wheel is 48 inches, how long is each spoke? Use 3.14 for Pi, and round your answer to the nearest hundredth. Need help with setting up the equation. Also confused about the "spikes"

OpenStudy (tylerd):

the spike will be the radius

OpenStudy (tylerd):

\[circumference=2\pi R\]

OpenStudy (anonymous):

Ok! Got it! so,

OpenStudy (tylerd):

in this case it says to use 3.14 for pi

OpenStudy (anonymous):

If would look something like this then:

OpenStudy (anonymous):

C=(2)(3.14)(48)?

OpenStudy (tylerd):

C=48

OpenStudy (anonymous):

where I take half of 48 for: C=(2)(3.14)(24) Am I correct? :)

OpenStudy (tylerd):

no

OpenStudy (anonymous):

Oh, messed up :(

OpenStudy (tylerd):